The following table shows prices and quantities in the hypothetical economy of Lowlands for two different items in two different years. Assume these are the only two items produced in this economy.

Prices and Output in a Two-Item Economy

Year 1 Year 2
Good Price Output Price Output
Apples $1.10 460 $1.30 520
Shirts 11.60 80 13.60 100

Instructions: Enter your answers as a whole number.

a. What is the value of nominal GDP for Lowlands in year 1?

     $

b. What is the value of nominal GDP for Lowlands in year 2?

     $   

c. Assume year 1 is used as the base year for this economy. What is the value of real GDP in year 2?

     $

d. In Lowlands, if nominal GDP were used instead of real GDP, the value of output in year 2 would be  (Click to select)  the same  understated  overstated  .

rev: 05_25_2018
Answer #1

a) Nominal GDP in year 1 = Price of apples (year 1) * quantity of apples (year 1) + price of shirts (year 1) * quantity of shirts (year 1) = 1.1*460 + 11.60*80 = 1434

b) Nominal GDP in year 2 = Price of apples (year 2) * quantity of apples (year 2) + price of shirts (year 2) * quantity of shirts (year 2) = 1.3*520 + 13.60*100 = 2036

c) Real GDP in year 2 = Price of apples (year 1) * quantity of apples (year 2) + price of shirts (year 1) * quantity of shirts (year 2) = 1.1*520 + 11.60*100 = 1732

d) It is overstated because real GDP is actually lower in year 2